# Provably Accurate Shapley Value Estimation via Leverage Score Sampling

### v1

@@ 6 Conclusion
We introduce Leverage SHAP, a principled alternative to Kernel SHAP, designed to provide provable accuracy guarantees with nearly linear model evaluations. The cornerstone of our approach is leverage score sampling, a powerful subsampling technique used in regression. To adapt this method for estimating Shapley values, we reformulate the standard Shapley value regression problem and analytically compute the leverage scores of this reformulation. Leverage SHAP efficiently uses these scores to produce provably accurate estimates of Shapley values. Our method enjoys strong theoretical guarantees, which we prove by modifying the standard leverage score analysis to incorporate the empirically-motivated paired sampling and sampling without replacement optimizations. Through extensive experiments on eight datasets, we demonstrate that our algorithm outperforms even the optimized version of Kernel SHAP, establishing Leverage SHAP as a valuable tool in the explainable AI toolkit.
@@ Appendix A Proof of Approximation Guarantee
Let $m=O(n\log(n/\delta)+n\log(1/\delta)/\epsilon)$. Algorithm 1 produces an estimate $\tilde{\bm{\phi}}$ such that, with probability $1-\delta$,
@@
Because of the connection between the constrained and unconstrained regression problems described in Lemma 3.1, the theorem implies the theoretical guarantees in Theorem 1.1. The time complexity bound in Theorem 1.1 follows from Lemma G.1 in Appendix G.
@@
$\|\mathbf{A\tilde{\bm{\phi}}-A\bm{\phi}}\|_{2}^{2}=\|\tilde{\mathbf{y}}-\mathbf{y}\|_{2}^{2}\leq 2\|\mathbf{U^{\top}S^{\top}SU}(\tilde{\mathbf{y}}-\mathbf{y})\|_{2}\leq\epsilon\|\mathbf{Uy-b}\|_{2}=\epsilon\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$ (38)
with probability $1-\delta$.
@@ Appendix B Proof of Approximation Corollary
Appendix C Additional Experiments: Performance by $\gamma$
The problem-specific term $\gamma=\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}/\|\mathbf{b}\|_{2}^{2}$ appears in the $\ell_{2}$-norm guarantee of Corollary 4.1.
In Figure 5, we explore experimentally whether $\gamma$ is an artifact of our analysis or a necessary parameter. We perform the experiment by explicitly building the matrix $\mathbf{A}$ (only possible for small $n$) and then choosing $\mathbf{b}$ as a linear combination of a vector in the column span of $\mathbf{A}$ and a vector not in the column span of $\mathbf{A}$. Then, we back out the corresponding set function $v$ using the definition of $\mathbf{b}$. We induce different values of $\gamma$ as we vary how much of $\mathbf{b}$ is a vector in the column span of $\mathbf{A}$. Because all three of the algorithms we consider perform worse as $\gamma$ increases, the experiment suggests that $\gamma$ is not an artifact of our analysis. The same trend appears for the ablated estimators, as shown in Figure 6. Encouragingly, $\gamma$ does not appear to be a relevant factor for the objective error, as shown in Figure 7.
@@ Appendix C Additional Experiments: Performance by $\gamma$
Appendix D Additional Experiments: $\ell_{2}$-norm error with Ablation
@@
Kernel SHAP
Mean 0.00261 0.0208 15.4 0.000139 0.00298 0.00324 0.0358 130.0
1st Quartile 5.69e-07 0.0031 3.71 1.48e-05 0.00166 0.00163 0.0106 33.5
2nd Quartile 9.52e-06 0.0103 8.19 3.86e-05 0.00249 0.00254 0.0221 53.6
3rd Quartile 0.00181 0.029 20.1 0.000145 0.00354 0.00436 0.0418 132.0
Optimized Kernel SHAP
Mean 3.28e-14 0.00248 2.33 1.81e-05 0.000739 0.000649 0.00551 21.8
1st Quartile 2.12e-14 0.000279 0.549 2.16e-06 0.00027 0.000187 0.000707 5.85
2nd Quartile 3.55e-14 0.00138 1.26 5.43e-06 0.000546 0.000385 0.0024 13.0
3rd Quartile 4.22e-14 0.0036 3.03 1.63e-05 0.00101 0.000964 0.00665 25.1
Leverage SHAP
Mean 3.28e-14 0.000186 0.63 5.21e-06 0.000458 0.000359 0.00385 14.7
1st Quartile 2.12e-14 1.91e-05 0.0631 6.3e-07 0.000139 9.51e-05 0.000333 3.6
2nd Quartile 3.55e-14 8.31e-05 0.328 2.33e-06 0.000376 0.000235 0.00149 8.9
3rd Quartile 4.22e-14 0.000231 0.769 7.09e-06 0.000617 0.000556 0.00401 15.3
Kernel SHAP Paired
Mean 0.000117 0.000539 1.13 7.61e-06 0.000682 0.000621 0.00589 24.6
1st Quartile 4.06e-14 4.15e-05 0.174 9.07e-07 0.000209 0.000148 0.000566 6.6
2nd Quartile 1.57e-09 0.00022 0.566 2.45e-06 0.000539 0.000389 0.00188 13.0
3rd Quartile 2.26e-06 0.000586 1.55 8.08e-06 0.000957 0.00097 0.00609 25.0
Leverage SHAP wo Bernoulli
Mean 0.000144 0.000478 1.05 8.07e-06 0.000445 0.000374 0.00293 13.6
1st Quartile 4.06e-14 3.46e-05 0.146 6.99e-07 0.000136 9.91e-05 0.0003 3.38
2nd Quartile 3.83e-10 0.000206 0.433 2.84e-06 0.000295 0.000271 0.00109 7.62
3rd Quartile 3.96e-06 0.000492 1.29 9.13e-06 0.00064 0.000562 0.00329 13.3
Leverage SHAP wo Bernoulli, Paired
Mean 0.0019 0.0165 13.3 0.000106 0.00202 0.00207 0.0207 73.1
1st Quartile 2.01e-07 0.00302 2.15 1.37e-05 0.000999 0.00099 0.00631 17.3
2nd Quartile 8.77e-06 0.00882 5.58 4.16e-05 0.00166 0.00172 0.0131 30.2
3rd Quartile 0.000348 0.0226 15.9 0.000104 0.00245 0.00278 0.0251 65.2
Table 2: Summary statistics of the $\ell_{2}$-norm error for every dataset and estimator. Except for ties, Leverage SHAP gives the best performance for small $n$. For large $n$, Leverage SHAP with replacement (without Bernoulli) sampling can mildly outperform.
Appendix E Additional Experiments: Linear Objective Error with Ablation
@@
Appendix F Equivalence
@@
Appendix G Sampling Without Replacement
Since there are an exponential number of rows, the naive method of considering each row independently will not work. Our approach is as follows: For each set size $s\in[\lfloor n/2\rfloor]$, we sample the number of samples that will be chosen. This is distributed as a Binomial distribution with $\binom{n}{s}$ trials and probability $\min(1,2c\binom{n}{s}^{-1})$. Then, because the probabilities are the same for all subsets of size $s$, we exploit this symmetry and sample random indices uniformly from $[\binom{n}{s}]$ without replacement. We index all subsets of size $s$ and then construct the subsets corresponding to the randomly sampled indices without enumerating all subsets. The Bernoulli sampling code appears in Algorithm 2 and the subset construction code appears in Algorithm 3, both deferred to Appendix G.
@@ Appendix G Sampling Without Replacement
Lemma G.1 (Time Complexity).
@@
Appendix H Constrained Regression
@@
$\nabla_{\lambda}\mathcal{L}(\mathbf{x}^{*},\lambda^{*})={\langle\mathbf{x}^{*}},\mathbf{1}\rangle-v(\mathbf{z})+v(\mathbf{0})=0.$ (60)
@@
$\langle\mathbf{1},{\mathbf{x}^{*}}\rangle=v(\mathbf{z})-v(\mathbf{0})=\mathbf{1}^{\top}(\mathbf{Z^{\top}Z})^{-1}\left(\mathbf{Z^{\top}b}-\frac{\lambda^{*}}{2}\mathbf{1}\right)$ (61)
@@
$\mathbf{x}^{*}=(\mathbf{Z^{\top}Z})^{-1}\left(\mathbf{Z^{\top}y}-\frac{\mathbf{1}^{\top}(\mathbf{Z^{\top}Z})^{-1}\mathbf{y}-v(\mathbf{1})+v(\mathbf{0})}{\mathbf{1}^{\top}(\mathbf{Z^{\top}Z})^{-1}\mathbf{1}}\mathbf{1}\right).$ (62)

### v2

@@ 6 Conclusion
We introduce Leverage SHAP, a principled alternative to Kernel SHAP, designed to provide provable accuracy guarantees with nearly linear model evaluations. The cornerstone of our approach is leverage score sampling, a powerful subsampling technique used in regression. To adapt this method for estimating Shapley values, we reformulate the standard Shapley value constrained regression problem and analytically compute the leverage scores of this reformulation. Leverage SHAP efficiently uses these scores to produce provably accurate estimates of Shapley values. Our method enjoys strong theoretical guarantees, which we prove by modifying the standard leverage score analysis to incorporate the empirically-motivated paired sampling and sampling without replacement optimizations. Through extensive experiments on eight datasets, we demonstrate that our algorithm outperforms even the optimized version of Kernel SHAP, establishing Leverage SHAP as a valuable tool in the explainable AI toolkit.
@@ Appendix A Proof of Approximation Guarantee
Let $m=O(n\log(\frac{n}{\delta})+n\frac{1}{\delta\epsilon})$. Algorithm 1 produces an estimate $\tilde{\bm{\phi}}$ such that, with probability $1-\delta$,
@@
Because of the connection between the constrained and unconstrained regression problems described in Lemma 3.1, the theorem implies the theoretical guarantees in Theorem 1.1. The time complexity bound in Theorem 1.1 follows from Lemma I.1 in Appendix I.
@@
$\|\mathbf{A\tilde{\bm{\phi}}-A\bm{\phi}}\|_{2}^{2}=\|\tilde{\mathbf{y}}-\mathbf{y}\|_{2}^{2}\leq 2\|\mathbf{U^{\top}S^{\top}SU}(\tilde{\mathbf{y}}-\mathbf{y})\|_{2}^{2}\leq\epsilon\|\mathbf{Uy-b}\|_{2}^{2}=\epsilon\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$ (38)
with probability $1-\delta$. ∎
@@ Appendix B Proof of Approximation Corollary
Appendix C Computational Hardness
When the Shapley value problem is viewed as an optimization problem, we provide a constant factor approximation in polynomial time (see e.g., Theorem 1.1).
On the other hand, suppose we wanted to, e.g., ensure that $\frac{1}{C}\phi_{i}\leq\tilde{\phi}_{i}\leq C\phi_{i}$ for all $i$. We claim that obtaining this goal for any constant $C$ is NP-hard in many settings. In particular, to ask about computational hardness, we also need to specify the input to the algorithm: i.e., is $v$ given as a polynomial size circuit, a polynomial size formula, or as a black-box, unit cost oracle? In all of these settings the problem is NP-hard to approximate. In particular, consider the case when $v$ either evaluates to 00 for all sets, or evaluates to 00 for all sets except for a single set $S$. In the first case, the Shapley values will all equal 00, whereas in the second case, the Shapley values for indices in $S$ are non-zero. As such, to obtain a multiplicative approximation, we need to determine if there is some set $S$ for which $v$ does not evaluate to 0. When $v$ is a black-box, finding such a set clearly requires $\Omega(2^{n})$ time (we can only enumerate all possible inputs). However, it is still hard when $v$ is given other forms. For example, if $v$ is a given as a circuit, then determining if there is an S for which $v(S)\neq 0$ is equivalent to the NP-complete circuit SAT problem.
This argument makes it clear why asking for a multiplicative approximation is not reasonable, and why instead we might care about an approximation in objective value. This is the case for many computational problems: for example, in the class of NP-hard optimization problems like set cover, we do not typically ask to approximate the optimal solution itself, but instead to find another solution with nearly the same objective value as the optimal solution (whether or not it is similar to that optimal solution or not).
Appendix D Noisy Access to the Set Function
In this section, we explore how the regression-based estimators perform given noisy access to the set function $v$. This setting is particularly relevant for Shapley values in explainable AI, since the set function may be an expectation that is approximated in practice. Figure 4 shows plots performance as noise is added to the set functions: Instead of observing $v(S)$, the estimators observe $v(S)+\zeta$ where $\zeta\sim\mathcal{N}(0,\sigma^{2})$ is normally distributed with standard deviation $\sigma$ (we set $\sigma=0,5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},\ldots,1$).
Appendix E Exploring $\gamma$
The problem-specific term $\gamma=\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}/\|\mathbf{b}\|_{2}^{2}$ appears in the $\ell_{2}$-norm guarantee of Corollary 4.1. In this section, we explore the distribution of $\gamma$, and the performance by $\gamma$. Because exactly computing $\gamma$ requires evaluating $\mathbf{b}$, we focus on datasets where $n\leq 16$.
Table 2 shows summary statistics of $\gamma$ over 100 runs (the randomness comes from the different models trained on the data). On every dataset, the third quartile of $\gamma$ is less than 2.
Table 2: Summary statistic of $\gamma$ over 100 runs.
Dataset $n$ 1st Quartile 2nd Quartile 3rd Quartile
IRIS 4 0.000234 0.266 1.03
California 8 0.158 0.298 0.449
Diabetes 10 0.174 0.321 0.513
Adult 12 0.180 0.395 0.703
In Figure 5, we explore experimentally whether $\gamma$ is an artifact of our analysis or a necessary parameter. We perform the experiment by explicitly building the matrix $\mathbf{A}$ (only possible for small $n$) and then choosing $\mathbf{b}$ as a linear combination of a vector in the column span of $\mathbf{A}$ and a vector not in the column span of $\mathbf{A}$. Then, we back out the corresponding set function $v$ using the definition of $\mathbf{b}$. We induce different values of $\gamma$ as we vary how much of $\mathbf{b}$ is a vector in the column span of $\mathbf{A}$. Because all three of the algorithms we consider perform worse as $\gamma$ increases, the experiment suggests that $\gamma$ is not an artifact of our analysis. The same trend appears for the ablated estimators, as shown in Figure 6. As our analysis suggests, $\gamma$ does not appear to be a relevant factor for the objective error, as shown in Figure 7.
@@ Appendix E Exploring $\gamma$
Appendix F Ablation Experiments with $\ell_{2}$-norm Error
Figures 8, 9, and 10 show ablation experiments where the error is measured with the $\ell_{2}$-norm error. We find that Leverage SHAP, or its ablated version without Bernoulli sampling, give the best performance.
@@ Appendix F Ablation Experiments with $\ell_{2}$-norm Error
Appendix G Ablation Experiments with Objective Error
Figures 11 and 12 show ablation experiments where the error is measured with the objective value naturally suggested by the regression formulation. We find that Leverage SHAP, or its ablated version without Bernoulli sampling, give the best performance.
@@ Appendix G Ablation Experiments with Objective Error
Appendix H Constrained and Unconstrained Equivalence
@@
Appendix I Efficiently Sampling Without Replacement
Since there are an exponential number of rows, the naive method of considering each row independently will not work. Our approach is as follows: For each set size $s\in[\lfloor n/2\rfloor]$, we sample the number of samples that will be chosen. This is distributed as a Binomial distribution with $\binom{n}{s}$ trials and probability $\min(1,2c\binom{n}{s}^{-1})$. Then, because the probabilities are the same for all subsets of size $s$, we exploit this symmetry and sample random indices uniformly from $[\binom{n}{s}]$ without replacement. We index all subsets of size $s$ and then construct the subsets corresponding to the randomly sampled indices without enumerating all subsets. The Bernoulli sampling code appears in Algorithm 2 and the subset construction code appears in Algorithm 3, both deferred to Appendix I.
@@ Appendix I Efficiently Sampling Without Replacement
Lemma I.1 (Time Complexity).
@@
Appendix J Constrained Regression
@@
$\nabla_{\lambda}\mathcal{L}(\mathbf{x}^{*},\lambda^{*})={\langle\mathbf{x}^{*}},\mathbf{1}\rangle-v(\mathbf{1})+v(\mathbf{0})=0.$ (60)
@@
$\langle\mathbf{1},{\mathbf{x}^{*}}\rangle=v(\mathbf{1})-v(\mathbf{0})=\mathbf{1}^{\top}(\mathbf{Z^{\top}Z})^{-1}\left(\mathbf{Z^{\top}y}-\frac{\lambda^{*}}{2}\mathbf{1}\right)$ (61)
@@
$\mathbf{x}^{*}=(\mathbf{Z^{\top}Z})^{-1}\left(\mathbf{Z^{\top}y}-\frac{\mathbf{1}^{\top}(\mathbf{Z^{\top}Z})^{-1}\mathbf{Z}^{\top}\mathbf{y}-v(\mathbf{1})+v(\mathbf{0})}{\mathbf{1}^{\top}(\mathbf{Z^{\top}Z})^{-1}\mathbf{1}}\mathbf{1}\right).$ (62)